Maude^{TM} License Agreement

Computer Science Laboratory SRI International Menlo Park CA 94025 USA

                   Email: maude-request@csl.sri.com
               www: http://www.csl.sri.com/maude.html

Phone: +1 (650) 859-4046 Fax: +1 (650) 859-2844

December 23, 1998

Please provide the information requested on pages 1, 4, and (if outside the USA) 5, sign, and return to Dr. Steven Eker at the address above.

There is no need to complete a new agreement if you already have a properly executed license for an earlier version of Maude.

Although not part of the agreement, we appreciate it if you will inform us at maude-bugs@csl.sri.com of any bugs that you discover, and notify us at maude-request@csl.sri.com of interesting applications you have made of Maude and of any papers or reports describing your use of the system.

SRI INTERNATIONAL END-USE LICENSE AGREEMENT

SOFTWARE PROGRAM: MAUDE^{TM}

GRANT

SRI International (“SRI”) hereby grants to a royalty-free, non-exclusive, non-transferable license to use its accompanying, above ref-

enter business name

erenced software program in executable form only (“Software”) and related documentation (“Documentation”) upon the terms and conditions set forth below:

You may: use the Software on any computers at your site; copy the Software for archival purposes, provided any copy must contain all of the original Software’s proprietary notices.

You may not: modify, translate, reverse engineer, decompile, disassemble (except to the extent applicable laws specifically prohibit such restriction), or create derivative works based on the Software or Documentation; copy the Software or Documentation (except for back-up purposes); rent, lease, transfer or otherwise transfer rights to the Software or Documentation; remove any proprietary notices or labels on the Software or Documentation; or permit others to use the Software except under the terms listed above.

TITLE

SRI retains title, ownership rights, and intellectual property rights in and to the Software and Documentation. The Software is protected by the copyright laws of the United States and international copyright treaties.

DISCLAIMER OF WARRANTY

Since this Software is a research tool still in the development stage and is provided free of charge, the Software and Documentation are provided on an “AS IS” basis, WITH-OUT WARRANTY OF ANY KIND, INCLUDING WITHOUT LIMITATION THE WAR-RANTIES OF MERCHANTABILITY OR FITNESS FOR ANY PARTICULAR PUR-POSE OR THAT THE USE OF THE SOFTWARE OR DOCUMENTATION WILL NOT INFRINGE ANY PATENTS, COPYRIGHTS, TRADEMARKS, OR OTHER RIGHTS. The entire risk as to the quality and performance of the Software and/or Documentation is borne by you. Should the Software and/or Documentation prove defective, you and not SRI assume the entire cost of any services and repair. This disclaimer of warranty constitutes an essential part of the agreement.

LIMITATION OF LIABILITY

UNDER NO CIRCUMSTANCES AND UNDER NO LEGAL THEORY, TORT, CON-TRACT, OR OTHERWISE, SHALL SRI, SRI’S AGENTS, OFFICERS, ASSISTANTS AND EMPLOYEES EITHER IN THEIR INDIVIDUAL CAPACITIES OR BY REASON OF THEIR RELATIONSHIP TO SRI AND SRI’S SUCCESSORS, BE HELD LIABLE TO YOU OR ANY OTHER PERSON FOR ANY INDIRECT, SPECIAL, INCIDENTAL, OR CONSEQUENTIAL DAMAGES OF ANY CHARACTER INCLUDING, WITHOUT LIMITATION, DAMAGES FOR LOSS OF GOODWILL, WORK STOPPAGE, COM-PUTER FAILURE OR MALFUNCTION, OR ANY AND ALL OTHER COMMERCIAL DAMAGES OR LOSSES. IN NO EVENT SHALL SRI BE LIABLE FOR COSTS OF PRO-CUREMENT OF SUBSTITUTE OR REPLACEMENT PRODUCTS OR SERVICES.

NO SUPPORT, NO UPDATES

This license does not grant you any right to any Software support, enhancements or updates.

TERMINATION

This license will terminate automatically if you fail to comply with the limitations described above. On termination, you must destroy all copies of the Software.

EXPORT CONTROLS

None of the Software or any underlying information or technology may be downloaded or otherwise exported or re exported (i) into (or to a national or resident of) Cuba, Iraq, Libya, Serbia, Montenegro, Hercegovina, Croatia, People’s Republic of China, any of the countries comprising the Commonwealth of Independent States (formerly the Soviet Union), North Korea, Iran, Syria or any other country to which the U.S. has embargoed goods; or (ii) to any person or entity listed as a Specially Designated National by the U.S. Treasury Department or the U.S. Commerce Department in its Table of Deny Orders. By downloading or using the Software, you are agreeing to the foregoing and you are representing and warranting that you are not located in, under the control of, or a national or resident of any such country or on any such list.

U.S. GOVERNMENT RESTRICTED RIGHTS

If the Software or any accompanying documentation is used or acquired by or on behalf of any unit, division or agency of the United States Government, this provision applies. The Software and any accompanying documentation is provided with RESTRICTED RIGHTS. The use, modification, reproduction, release, display, duplication or disclosure thereof by or on behalf of any unit, division or agency of the Government is subject to the restrictions set forth in subdivision (c)(1) of the Commercial Computer Software-Restricted Rights clause at 48 CFR 52.227-19 and the restrictions set forth in the Rights in Technical Data-Non-commercial Items clause set forth in 48 CFR 252.227-7013. The contractor/manufacturer of the Software and accompanying documentation is SRI International, 333 Ravenswood Avenue, Menlo Park, California 94025, USA.

MISCELLANEOUS

This agreement represents the complete agreement concerning this license between the parties and supersedes all prior agreement and representations between them. It may be amended only by a writing executed by both parties. If any provision of this Agreement is held to be unenforceable for any reason, such provision shall be reformed only to the extent necessary to make it enforceable. This agreement shall be governed by and construed in accordance with the laws of the State of California, without regard to the conflicts of law principles thereof, and shall not be governed by the United Nations convention on contracts for the international sale of goods.

SIGNATURES

The licensed material is provided for the use of LICENSEE at the following

SITE/FACILITY only:  Enter location where Maude will be used

Facility: Address:

Enter name and address of person for correspondence

Name: Address:

(leave address blank if same as above)

Phone: Fax: Email:

In the spaces below on the right, enter the name of your business or institution. The person signing below the business name should be someone who has the legal authority to sign for your business or institution. That person’s position of authority (title) in the business or institution should be entered on the next line. Return the completed license to SRI at the address on the cover; SRI will countersign on the left below and return to you at the address you enter above.

LICENSOR: SRI International

By

Title

Date LICENSEE:

By

Title

Date END-USE STATEMENT

enter business name enter signature enter business title

Use of Maude outside the USA may be subject to US Government export licensing requirements. The requirements are dependent on the use to be made of Maude. Overseas users must therefore submit with their completed license a statement of the end-use to which they plan to put the software.

If it applies to your use of Maude, we suggest you use the following form of words:

Maude will be used to conduct basic and applied research. It is intended that the results of the research will be made generally available, and published should they merit it.

If your use of Maude will concern something other than research, or if the results will be restricted for proprietary or national security reasons, then please provide explicit details in your end-use statement.

Please write or type your end-use statement below.